OpenStudy (anonymous):

Calculate f ′(t). f (t) = (e^(12t)i + 20tj) • (i + 2j) (Note: 12t is a power to e)

OpenStudy (amistre64):

looks like the function itself is a dot product of 2 vectors

OpenStudy (amistre64):

id try to apply a product rule for vectors :)

OpenStudy (anonymous):

Alright, I'll try to see if that works.

OpenStudy (amistre64):

f(t) = r(t) l(t) f'(t) = r(t) l(t)' + r(t)' l(t) r(t) = <e^(12t),20t> r(t)'= <12 e^(12t), 20> l(t) = <1,2> l(t)'=<0,0> if im remembering this correctly

OpenStudy (amistre64):

f(t)' = <e^(12t),20t> + <12 e^(12t), 20> < 0 , 0 > < 1 , 2 > -------------------------------- 0+0 + 12e^(12t) + 40

OpenStudy (anonymous):

Hey, I was solving it out and I got the same answer as yours, but I did not use your method. I've used the method that my professor gave and that works as well. Thanks for helping me. :)

OpenStudy (amistre64):

yep, :) good luck